# Localized Band Gap Resonances in Defect-embedded Monoatomic Chains

### v1

@@ Abstract
Historically, structural defects in crystalline material lattice structures have been viewed as a liability, causing the material behavior to deviate from the expected norm. However, the gradual improvement in understanding the role of defects has overturned this view, encouraging positive outlooks, e.g., deliberately engineering defects in structures, that have laid the foundation for modern applications, e.g., silicon chip manufacturing, battery engineering, fuel cell engineering. In recent years, acoustic metamaterials (AMMs), which have emerged as a platform for customizing acoustic behavior, have also embraced this shift in viewpoint, enabling design propositions involving precisely embedded defects in the AMM architecture that may be relevant for applications, e.g., energy harvesting, waveguiding, flow control. The fundamental utility of defects stems from their ability to robustly localize energy at a prescribed location with desirable spatio-temporal characteristics. However, the bulk of existing defect-based AMM studies primarily present a qualitative numerical analysis, prompting the need for an analytical approach quantifying the defect eigenmode characteristics, e.g., frequency, mode shape, resonance velocity amplitude envelope to precisely engineer defects in AMMs. This article presents a modified analytical method to determine the eigenmode characteristics of a defect-embedded grounded monoatomic chain (DMoA) and demonstrates the benefits of defect modes through two DMoA-based AMMs that can accommodate – (i) a virtual ground, and (ii) achieve customized acoustic interaction and absorption. The proposed strategy can be readily extended to more elaborate AMMs and augments the design space for defect-based AMMs.
Keywords
Metamaterials; bandgap resonances; defect modes; energy localization
I Introduction
@@
Simultaneously, in the field of mechanics and materials, the advent of acoustic metamaterials (AMMs), primarily featuring periodic, cleverly architected mesoscale building blocks (i.e., unit cells) have offered an invaluable capability to mimic the ordered atomic architectures of natural materials at a more accessible scale and tailor the macroscale material behavior. AMMs have featured diverse unit cell designs, leveraging active elements[9, 10], geometric instabilities[11, 12], multistability[13, 14], and kinematic amplification mechanisms[15, 16], eliciting peculiar material behavior, e.g., auxeticity[17], non-reciprocity[10], directional wave-guiding[18]. Though AMM research is relatively recent, the community has shared identical views with the larger materials research community about the presence of structural or material defects[19]. Similar to the prevalent perception of defects, this view has continuously evolved and, in recent years, has led to the exploration of material architectures with precisely embedded defects to robustly localize energy at desirable locations and spatio-temporal scales. Multiple strategies[20, 21, 22, 23, 24] exist in the literature, to introduce defect modes in a variety of material architectures, e.g., Kaina et al.[20] experimentally demonstrated a waveguide allowing wave propagation at reduced group velocities by altering the lattice spacing of the periodic sub-wavelength resonant structures, Li et al.[21] induced a defect mode in a soft bilayer by triggering a wrinkling instability at a prescribed location via strain-engineering, Xu-Feng et al.[22] experimentally demonstrated high energy localization and efficient energy harvesting by inducing geometric defects in a phononic crystal (PnC) beam. These proposals primarily present a qualitative analysis of the defect resonance dynamics via numerical simulations, parameter studies, supercell analysis, and experiments.
@@
Consequently, subsequent studies have explored analytical methods[25, 26, 27, 28, 29, 30, 31, 32, 33, 34] to characterize the defect mode frequency in terms of the defect parameters, e.g., geometric parameters, material properties. Though these approaches have aided the development of more informed material design strategies, the frequency represents only a subset of the relevant dynamic characteristics of defect modes. Additional modal characteristics, e.g., mode shape, localized velocity (displacement) at the defect location, can also play a significant role in determining the effectiveness of defect-based AMMs and augment their functionality for wave-guiding[35], energy harvesting[36, 37], flow control[38] applications, motivating the need for a more comprehensive characterization of defect modes. Addressing this demand, we present a modified analytical approach to quantify three critical characteristics of defect modes, i.e., frequency, mode shape, and resonance velocity amplitude envelope, in a defect-embedded grounded monoatomic chain.
@@
The remainder of the article is organized as follows – Sec. II provides a comprehensive analysis of the acoustic characteristics of the emergent defect resonances in a defect-embedded grounded monoatomic chain. Sec. III presents two defect-embedded AMMs described in Sec. II, demonstrating the benefits of defect modes. Finally, Sec. IV presents a summary and the key takeaways from the article.
II Model Description
@@
Figure 1: Defect-embedded grounded monoatomic chain (DMoA). (a) Mass-spring model for a defect-free, uniform grounded monoatomic chain (MoA), dispersion curves with pass bands (PBs), band gaps (BGs), and eigenfrequencies of finite ($\mathrm{N}=11$) edge and bulk DMoAs. Mass-spring models and frequency response functions (FRFs) of a finite (b) Edge DMoA, and (c) Bulk DMoA.
@@
Consider a periodic mass-spring model of a 1D grounded monoatomic chain (MoA) as shown in Fig. 1a with mass, $m$, interaction stiffness, $k$ and onsite grounding stiffness, $k_{\mathrm{g}}$. The non-dimensional equation of motion for the $j^{\mathrm{th}}$ mass in an infinite MoA is given by,
@@
where, $v_{j}$ is the axial displacement of the $j^{\mathrm{th}}$ mass, $\bar{t}=\omega_{0}t$ is the non-dimensional time, $\omega_{0}=\sqrt{k/m}$ is the reference temporal frequency and $k_{\mathrm{g,r}}=k_{\mathrm{g}}/k$ is the grounding stiffness ratio. Consequently, the dispersion relation for the MoA, assuming a plane wave ansatz of the form $v_{j}=\tilde{v}_{j}e^{i(\kappa x_{j}-\Omega\bar{t})}\ \forall\ j=1,2,\cdots,\mathrm{N}\ \mathrm{and}\ x_{j}=j-1$, is calculated as[39, 40]:
@@
where, $\Omega$ is the non-dimensional temporal frequency, and $\kappa=\kappa_{R}+i\kappa_{I}\ \forall\ \kappa_{R},\kappa_{I}\in\mathbb{R}$, is the non-dimensional complex wavenumber. The dispersion curves (Eq. (2), $m=1$, $k=1$, $k_{\mathrm{g,r}}=2$) and the natural eigenfrequencies, $\Omega_{j}$ of a finite MoA with $\mathrm{N}$ degrees of freedom and free boundaries are plotted in Fig. 1a, highlighting the band gap (BG, $\kappa_{R}=0\ \mathrm{or}\ \pi,\kappa_{I}\neq 0$) and the pass band (PB, $\kappa_{R}\neq 0,\kappa_{I}=0$) frequency ranges. The grounding elastic springs are the primary source of the bounded BG between $\Omega=0$ and a finite $\Omega=\sqrt{k_{\mathrm{g,r}}}$. The grounding springs resist the deformation of masses with respect to a fixed ground and participate in all the eigenmodes describing the MoA dynamics. Consequently, the PB range and specifically the lower edge of the PB, $\Omega(\kappa=0)=\sqrt{k_{\mathrm{g,r}}}$ is directly linked to the elasticity of these grounding elements, opening up a BG below this modal frequency.
@@
When the periodicity of a finite MoA is altered, by introducing a change in one or more material properties (defect), i.e., $m_{\mathrm{r}}=m_{\mathrm{d}}/m$, $k_{\mathrm{r}}=k_{\mathrm{d}}/k$, or $k_{\mathrm{gd,r}}=k_{\mathrm{gd}}/k$ at a specified location within the MoA, the eigenfrequency (or multiple eigenfrequencies, see Sec. A2) associated with the lower (upper) edge of the PB migrates to the lower (upper) BG, creating a BG resonance at $\Omega_{1}=\Omega_{\mathrm{d}}<\sqrt{k_{\mathrm{g,r}}}$ ($\Omega_{\mathrm{N}}=\Omega_{\mathrm{d}}>\sqrt{4+k_{\mathrm{g,r}}}$). Note that a BG resonance can exist even in an ungrounded mass-spring chain[30, 28, 41] however, a grounded chain ($k_{\mathrm{g,r}}\neq 0$) provides an opportunity to engineer a BG resonance in the lower BG, $\Omega_{\mathrm{d}}\in[0,\sqrt{k_{\mathrm{g,r}}})$, that manifests as the first eigenmode of the altered MoA. The existence of a BG resonance as the principal, low-frequency eigenmode may be beneficial in ensuring their effective participation in the system dynamics, facilitating DMoA-based AMMs for practical applications that can involve complex operational requirements, e.g., response to low-frequency broadband external forces[41], amplitude-phase requirements[38]. We refer to the altered MoA as a defect-embedded grounded monoatomic chain (DMoA) and the resultant BG resonance as a defect mode for brevity.
@@
The DMoA can be explored in two fundamentally distinct configurations – (i) an edge DMoA where material properties at one of the boundaries of the finite MoA are altered and, (ii) a bulk DMoA[32] where material properties in the bulk deviate from the periodic properties of the MoA. The natural frequencies of these DMoAs are obtained by solving the matrix eigenvalue problem, $\mathbf{K}_{\mathrm{d}}\hat{\mathbf{V}}=\mathbf{0}$ (see Sec. A1), with $m_{\mathrm{r}}=k_{\mathrm{r}}=1$ and $k_{\mathrm{gd,r}}=0$ and, overlayed onto the dispersion plot in Fig. 1a, indicating the presence of a defect mode (BG resonance) in both cases albeit, with distinct frequencies. In addition, the frequency response functions (FRFs) of the DMoAs, evaluated by applying an external force and monitoring the displacement at the defect mass are plotted in Fig. 1b,c. In either case, the DMoA exhibits relatively higher FRF amplitudes at the defect frequency, $\Omega_{\mathrm{d}}$ compared to other eigenfrequencies, indicating energy localization at the defect mass owing to the manifestation of an exponentially decaying eigenmode shape away from the defect location (see insets of Fig. 1b,c). The remainder of the article uses – $m=1$, $k=1$, $k_{\mathrm{g,r}}=2$, $m_{\mathrm{r}}=1$, $k_{\mathrm{r}}=1$ and $k_{\mathrm{gd,r}}=0$ unless explicitly stated otherwise.
@@
The defect frequency, $\Omega_{\mathrm{d}}$, and the exponential decay rate, $\kappa_{I}$ comprehensively describe the contribution of the material defect in altering the base MoA dynamics. In addition, although rarely considered in prior studies, we quantify and describe a third critical parameter – the velocity amplitude envelope at resonance, $\dot{\mathrm{V}}_{\mathrm{E}}$ that can promote defect-based AMMs for novel applications, e.g., energy harvesting[36, 37], flow control via fluid-metamaterial interactions[38, 42], where response amplitude of the AMM can play a major role in determining operational efficiency. In the following sections, we elaborate on a modified analytical approach to quantify these critical characteristics of the defect dynamics of the DMoA – $\{\Omega_{\mathrm{d}},\kappa_{I},\dot{\mathrm{V}}_{\mathrm{E}}\}$. The proposed approach can be readily extended to general boundary conditions[28, 30] and mass-spring AMM models comprising unit cells with more degrees of freedom, e.g., defect-embedded diatomic[26, 28] and polyatomic[29] chains.
II.1 Defect eigenfrequency and eigenmode shape
@@
Figure 2: Iso-curves of the defect eigenmode characteristics. (Center) 3D plot of the iso-surfaces of defect frequency, $\Omega_{\mathrm{d}}=\{0.4,0.6,0.8,1.0,1.2\}$ as a function of material properties, $m_{\mathrm{r}}$ (log-scale) and $k_{\mathrm{gd,r}}$ and, the size of the base MoA, $\mathrm{N}$. (a) Contour depicting a 2D slice of the 3D central figure, taken at $m_{\mathrm{r}}=1$, (b) Contour depicting a 2D slice of the 3D central figure, taken at $k_{\mathrm{gd,r}}=0$, (c-i) Contour depicting a 2D slice of the 3D central figure, taken at $\mathrm{N}=21$, (c-ii) Contour plotting the analytical solution for defect frequency, $\Omega_{\mathrm{d}}$ as a function of $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$, (c-iii) Numerical and (c-iv) analytical contours of the defect eigen mode shape, i.e., decay rate, $\kappa_{I}$ as a function of $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$. The $\{\square,\mathbf{O},\mathbf{\times},\bullet\}$-markers indicate DMoA configurations investigated in Fig. 3.
@@
Consider the edge DMoA model illustrated in Fig. 1b. The edge defect resembles a local resonator appended to the boundary of a finite, homogeneous elastic foundation (base MoA). The presence of the edge defect creates a localized defect mode in the altered material configuration that can be modeled as a reduced order edge DMoA illustrated in Fig. 1b. The BG eigenfrequencies of a finite edge DMoA are numerically calculated and plotted as iso-defect frequency surfaces, i.e., $\Omega_{\mathrm{d}}=\mathrm{const.}$, in the central 3D-plot in Fig. 2 for varying degrees of freedom, $\mathrm{N}$, and material properties at the edge defect, i.e., $k_{\mathrm{gd,r}}$ and $m_{\mathrm{r}}$, to understand the qualitative dependence of the defect frequency, $\Omega_{\mathrm{d}}$ on these parameters. Fig. 2a presents a 2D slice of the central plot taken at $m_{\mathrm{r}}=1$. Consistent with the qualitative relation between frequency and stiffness, the magnitudes of the iso-defect frequency curves on the contour plot indicate that, $\Omega_{\mathrm{d}}$ increases as $k_{\mathrm{gd,r}}$ increases from $0\longrightarrow k_{\mathrm{g,r}}$. Similarly, in Fig. 2b (2D slice taken at $k_{\mathrm{gd,r}}=0$) the magnitude of iso-defect frequency curves decrease with increasing $m_{\mathrm{r}}$ ($>1$), as anticipated qualitatively. In addition, we notice that the iso-defect frequency curves in both Fig. 2a,b tend to asymptote to a constant value of $k_{\mathrm{gd,r}}$ and $m_{\mathrm{r}}$, respectively, at large $\mathrm{N}$. This observation provides valuable insight for our analytical model, where we can arbitrarily choose a large DMoA size ($\mathrm{N}$) to obtain an accurate prediction for $\Omega_{\mathrm{d}}$, independent of the actual size of the finite DMoA.
@@
Finally, Fig. 2(c-i,iii) plots the $\{\Omega_{\mathrm{d}},\kappa_{I}\}$, calculated from matrix eigen-analysis of a finite DMoA ($\mathrm{N}=21$) as a function of $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$. A closer look at the matrix eigenvalue problem reveals a tridiagonal matrix form for the dynamic stiffness, $\mathbf{K}_{\mathrm{d}}$ (Sec. A1) that is consistent with the general form of a perturbed tridiagonal Toeplitz matrix[43, 28]. The characteristic equation for $\mathbf{K}_{\mathrm{d}}$ can be written in the implicit form:
@@
$f(\Omega)=\frac{\sin\left([n+1]\left[\cos^{-1}\left(\frac{d^{2}}{2a^{2}}-1\right)\right]\right)}{\sin\left(n\left[\cos^{-1}\left(\frac{d^{2}}{2a^{2}}-1\right)\right]\right)},$ (4)
@@
$d=2+k_{\mathrm{g,r}}-\Omega^{2},a=-1,\beta=-1,$ (5)

### v2

# A Quantitative Study of Energy Localization Characteristics in Defect-embedded Phononic Crystals

@@ Abstract
Phononic crystals (PnCs) are periodic engineered media that can customize the spatio-temporal characteristics of mechanical energy propagation. PnCs that additionally leverage precisely embedded defects can achieve robust energy localization with desirable spatio-temporal characteristics, opening avenues for critical engineering applications, e.g., energy harvesting, waveguiding, and fluid flow control. Numerous studies have qualitatively explored the localized dynamics via simulations and experiments, investigating the defect resonance frequency as the primary feature. However, the frequency represents only a subset of the relevant characteristics and a systematic approach to quantify the full scope of the defect dynamics remains elusive. This article establishes the frequency, mode shape, and localized velocity (or displacement) amplitude envelope as three significant factors governing the defect resonance dynamics, and quantitatively examines these characteristics using a modified version of the perturbed tridiagonal n-Toeplitz method. The proposed method accurately estimates the resonance characteristics in 1D and 2D defect-embedded PnC lattices with single and multiple defects and elucidates the effects of damping. The method is used to highlight how the key characteristics of defect modes depend on system parameters. Finally, we demonstrate the benefits of defect modes through two defect-based PnCs that can accommodate – (i) a virtual ground, and (ii) achieve customized acoustic interaction and absorption, and use the proposed method to analyze these scenarios. The proposed strategy can be readily extended to more elaborate PnCs and augments the design space for defect-based PnCs.
keywords:
Phononic crystals , bandgap resonances , defect modes , energy localization
[…]
[inst1]organization=The Grainger College of Engineering, Mechanical Science and Engineering, addressline=University of Illinois Urbana-Champaign, postcode=Illinois, 61801, city=Urbana, country=USA
1 Introduction
@@
Simultaneously, in the field of mechanics and materials, the advent of phononic crystals (PnCs) and metamaterials, primarily featuring periodic, cleverly architected mesoscale building blocks (i.e., unit cells) have offered an invaluable capability to mimic the ordered atomic architectures of natural materials at a more accessible scale and tailor the macroscale material behavior. PnCs and metamaterials have featured diverse unit cell designs, leveraging active elements[9, 10], geometric instabilities[11, 12], multistability[13, 14], and kinematic amplification mechanisms[15, 16], eliciting peculiar material behavior, e.g., auxeticity[17], non-reciprocity[10], directional wave-guiding[18]. Though this research is relatively recent, the community has shared identical views with the larger materials research community about the presence of structural or material defects[19]. Similar to the prevalent perception of defects, this view has continuously evolved and, in recent years, has led to the exploration of material architectures with precisely embedded defects to robustly localize energy at desirable locations and spatio-temporal scales. Multiple strategies[20, 21, 22, 23, 24] exist in the literature, to introduce defect modes in a variety of material architectures, e.g., Kaina et al.[20] experimentally demonstrated a waveguide allowing wave propagation at reduced group velocities by altering the lattice spacing of the periodic sub-wavelength resonant structures, Li et al.[21] induced a defect mode in a soft bilayer by triggering a wrinkling instability at a prescribed location via strain-engineering, Xu-Feng et al.[22] experimentally demonstrated high energy localization and efficient energy harvesting by inducing geometric defects in a PnC beam. These proposals primarily present a qualitative analysis of the defect resonance dynamics via numerical simulations, parameter studies, supercell analysis, and experiments.
@@
Consequently, subsequent studies have explored analytical methods[25, 26, 27, 28, 29, 30, 31, 32, 33, 34] to characterize the defect mode frequency in terms of the defect parameters, e.g., geometric parameters, material properties. Though these approaches have aided the development of more informed material design strategies, the frequency represents only a subset of the relevant dynamic characteristics of defect modes. Additional modal characteristics, e.g., mode shape, localized velocity (or displacement) amplitude envelope at the defect location, can also play a significant role in determining the effectiveness of defect-based PnCs and augment their functionality for wave-guiding[35], energy harvesting[36, 37], flow control[38] applications, motivating the need for a more comprehensive characterization of defect modes. Addressing this demand, we present a modified analytical approach to quantify three critical characteristics of defect modes, i.e., frequency, mode shape, and resonance velocity amplitude envelope, in a defect-embedded grounded monoatomic phononic crystal.
@@
The remainder of the article is organized as follows – Sec. 2 provides a comprehensive analysis of the acoustic characteristics of the emergent bandgap resonances in a defect-embedded grounded monoatomic phononic crystal. Sec. 3 presents two defect-embedded PnCs, demonstrating the benefits of defect modes. Finally, Sec. 4 presents a summary and the key takeaways from the article.
2 Model Description
@@
Figure 1: DMoA PnC. (a-i) A mass-spring MoA PnC model, dispersion curves with pass bands (PBs), band gaps (BGs), and eigenfrequencies of finite ($\mathrm{N}=11$) edge and bulk DMoA PnCs. (a-ii) A 3D structural model consisting of an array of aluminum plates ($m$), connected to their nearest neighbors using silicone beams ($k$) and individually grounded using Veroclear clips ($k_{\mathrm{g}}$) whose axial dynamics are represented by the MoA PnC model. (i) Mass-spring DMoA PnC models and frequency response functions (FRFs) obtained from 3D FEM studies of a finite defect-embedded 3D structure, (ii) defect eigenmode shapes featuring - (b) an edge defect, and (c) a bulk defect.
@@
Consider a composite 3D structure shown in Fig. 1(a-ii). The axial dynamics of the 3D structure can be approximated as a periodic lumped mass-spring grounded monoatomic phononic crystal (MoA PnC) model (Fig. 1(a-i)) with mass, $m$, interaction stiffness, $k$, and onsite grounding stiffness, $k_{\mathrm{g}}$ (see Sec. A7). The non-dimensional equation of motion for the $j^{\mathrm{th}}$ mass in an infinite MoA PnC is given by,
@@
where, $v_{j}$ is the axial displacement of the $j^{\mathrm{th}}$ mass, $\bar{t}=\omega_{0}t$ is the non-dimensional time, $\omega_{0}=\sqrt{k/m}$ is the reference temporal frequency and $k_{\mathrm{g,r}}=k_{\mathrm{g}}/k$ is the grounding stiffness ratio. Consequently, the dispersion relation for the MoA PnC, assuming a plane wave ansatz of the form $v_{j}=\tilde{v}_{j}e^{i(\kappa x_{j}-\Omega\bar{t})}\ \forall\ j=1,2,\cdots,\mathrm{N}\ \mathrm{and}\ x_{j}=j-1$, is calculated as[39, 40]:
@@
where, $\Omega$ is the non-dimensional temporal frequency, and $\kappa=\kappa_{R}+i\kappa_{I}\ \forall\ \kappa_{R},\kappa_{I}\in\mathbb{R}$, is the non-dimensional complex wavenumber. The dispersion curves (Eq. (2), $k_{\mathrm{g,r}}=2$) and the natural eigenfrequencies, $\Omega_{j}$ of a finite MoA PnC with $\mathrm{N}$ degrees of freedom and free boundaries are plotted in Fig. 1a, highlighting the band gap (BG, $\kappa_{R}=0\ \mathrm{or}\ \pi,\kappa_{I}\neq 0$) and the pass band (PB, $\kappa_{R}\neq 0,\kappa_{I}=0$) frequency ranges. The grounding elastic springs are the primary source of the bounded BG between $\Omega=0$ and a finite $\Omega=\sqrt{k_{\mathrm{g,r}}}$. The grounding springs resist the deformation of masses with respect to a fixed ground and participate in all the eigenmodes describing the MoA PnC dynamics. Consequently, the PB range and specifically the lower edge of the PB, $\Omega(\kappa=0)=\sqrt{k_{\mathrm{g,r}}}$ is directly linked to the elasticity of these grounding elements, opening up a BG below this modal frequency.
@@
When the periodicity of a finite MoA PnCs is altered by introducing a change in one or more material properties (defect), i.e., $m_{\mathrm{r}}=m_{\mathrm{d}}/m$, $k_{\mathrm{r}}=k_{\mathrm{d}}/k$, or $k_{\mathrm{gd,r}}=k_{\mathrm{gd}}/k$ at a specified location within the MoA PnC, the eigenfrequency (or multiple eigenfrequencies, see Sec. A2) associated with the lower (upper) edge of the PB migrates to the lower (upper) BG, creating a BG resonance at $\Omega_{1}=\Omega_{\mathrm{d}}<\sqrt{k_{\mathrm{g,r}}}$ ($\Omega_{\mathrm{N}}=\Omega_{\mathrm{d}}>\sqrt{4+k_{\mathrm{g,r}}}$). Note that a BG resonance can exist even in an ungrounded mass-spring PnCs[30, 28, 41] however, a grounded PnC ($k_{\mathrm{g,r}}\neq 0$) provides an opportunity to engineer a BG resonance in the lower BG, $\Omega_{\mathrm{d}}\in[0,\sqrt{k_{\mathrm{g,r}}})$, that manifests as the first eigenmode of the altered MoA PnC. The existence of a BG resonance as the principal, low-frequency eigenmode may be beneficial in ensuring their effective participation in the system dynamics, facilitating defect-based PnCs for practical applications that can involve complex operational requirements, e.g., response to low-frequency broadband external forces[41], amplitude-phase requirements[38]. We refer to the altered MoA PnC as a defect-embedded grounded monoatomic phononic crystal (DMoA PnC) and the resultant BG resonance as a defect mode for brevity.
@@
The DMoA PnCs can be explored in two fundamentally distinct configurations – (i) an edge DMoA PnC where material properties at one of the boundaries of the finite MoA PnC are altered and, (ii) a bulk DMoA PnC[32] where material properties in the bulk deviate from the periodic properties of the MoA PnC. The natural frequencies of these DMoA PnCs are obtained by solving the matrix eigenvalue problem, $\mathbf{K}_{\mathrm{d}}\hat{\mathbf{V}}=\mathbf{0}$ (see Sec. A1), with $m_{\mathrm{r}}=k_{\mathrm{r}}=1$ and $k_{\mathrm{gd,r}}=0$ and, overlaid onto the dispersion plot in Fig. 1a, indicating the presence of a defect mode (BG resonance) in both cases albeit, with distinct frequencies.
The frequency response functions (FRFs) of the 3D structures with edge and bulk defects, plotted in Fig. 1(b,c-i) are calculated by applying a uniform boundary load in the axial (z-) direction to the defect plate and performing a 3D FEM frequency analysis in COMSOL. Though modes of other polarizations (beam bending modes, see Fig. A5) are faintly excited in the edge defect-embedded structure ($\Omega\approx 2.9$), the FRF resonance peaks primarily denote the axial resonance modes of the structures, showing an excellent correlation with the eigenfrequencies calculated from the edge and bulk DMoA PnC mass-spring models. In addition, the FRF captures all $\mathrm{N}$ axial eigenmodes and only $(\mathrm{N}+1)/2$ axial eigenmodes, respectively, in the edge and bulk defect cases as the excitation location is at the structural boundary in the former as compared to the material bulk in the latter.
The 3D structures manifest higher FRF amplitudes at the defect resonance frequency, $\Omega_{\mathrm{d}}$ compared to other PB eigenfrequencies. A reduction in amplitudes between the magenta ($|\mathrm{A}|_{j=d}$) and black curves ($|\mathrm{A}|_{j=\mathrm{N}}$) in Fig. 1(b,c-i) at $\Omega=\Omega_{\mathrm{d}}$ also indicates a strong localization of energy at the defect owing to an exponentially decaying eigenmode shape away from the defect location (see Fig. 1(b,c-ii)). Since the axial modes of the 3D structures subjected to uniform axial loads are accurately captured by the reduced order model, the remainder of the article analytically and numerically investigates the DMoA PnC mass-spring model with material parameters – $k_{\mathrm{g,r}}=2$, $m_{\mathrm{r}}=1$, $k_{\mathrm{r}}=1$ and $k_{\mathrm{gd,r}}=0$ unless explicitly stated otherwise.
@@
The defect frequency, $\Omega_{\mathrm{d}}$, and the exponential decay rate, $\kappa_{I}$ comprehensively describe the contribution of the material defect in altering the base MoA PnC dynamics. In addition, although rarely considered in prior studies, we quantify and describe a third critical parameter – the velocity amplitude envelope at resonance, $\dot{\mathrm{V}}_{\mathrm{E}}$ that can promote defect-based PnCs for novel applications, e.g., energy harvesting[36, 37], fluid flow control[38, 42], where response amplitude of the PnC can play a major role in determining operational efficiency. In the following sections, we elaborate on a modified analytical approach to quantify these critical characteristics of the defect dynamics of the DMoA PnC – $\{\Omega_{\mathrm{d}},\kappa_{I},\dot{\mathrm{V}}_{\mathrm{E}}\}$. The proposed approach can be readily extended to general boundary conditions[28, 30] and mass-spring PnC models comprising unit cells with more degrees of freedom, e.g., defect-embedded diatomic[26, 28] and polyatomic[29] phononic crystals.
2.1 Defect eigenfrequency and eigenmode shape
@@
Figure 2: Iso-curves of the defect eigenmode characteristics. (Center) 3D plot of the iso-surfaces of defect frequency, $\Omega_{\mathrm{d}}=\{0.4,0.6,0.8,1.0,1.2\}$ as a function of material properties, $m_{\mathrm{r}}$ (log-scale) and $k_{\mathrm{gd,r}}$ and, the size of the base MoA PnC, $\mathrm{N}$. (a) Contour depicting a 2D slice of the 3D central figure, taken at $m_{\mathrm{r}}=1$, (b) Contour depicting a 2D slice of the 3D central figure, taken at $k_{\mathrm{gd,r}}=0$, (c-i) Contour depicting a 2D slice of the 3D central figure, taken at $\mathrm{N}=21$, (c-ii) Contour plotting the analytical solution for defect frequency, $\Omega_{\mathrm{d}}$ as a function of $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$, (c-iii) Numerical and (c-iv) analytical contours of the defect eigen mode shape, i.e., decay rate, $\kappa_{I}$ as a function of $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$. The $\{\square,\mathbf{O},\mathbf{\times},\bullet\}$-markers indicate DMoA PnC configurations investigated in Fig. 3.
@@
Consider the edge DMoA PnC model illustrated in Fig. 1b. The edge defect resembles a local resonator appended to the boundary of a finite, homogeneous elastic foundation (base MoA PnC). The presence of the edge defect creates a localized defect mode in the altered material configuration that can be modeled as a reduced order edge DMoA PnC illustrated in Fig. 1b. The BG eigenfrequencies of a finite edge DMoA PnC are numerically calculated and plotted as iso-defect frequency surfaces, i.e., $\Omega_{\mathrm{d}}=\mathrm{const.}$, in the central 3D-plot in Fig. 2 for varying degrees of freedom, $\mathrm{N}$, and material properties at the edge defect, i.e., $k_{\mathrm{gd,r}}$ and $m_{\mathrm{r}}$, to understand the qualitative dependence of the defect frequency, $\Omega_{\mathrm{d}}$ on these parameters. Fig. 2a presents a 2D slice of the central plot taken at $m_{\mathrm{r}}=1$. Consistent with the qualitative relation between frequency and stiffness, the magnitudes of the iso-defect frequency curves on the contour plot indicate that, $\Omega_{\mathrm{d}}$ increases as $k_{\mathrm{gd,r}}$ increases from $0\longrightarrow k_{\mathrm{g,r}}$. Similarly, in Fig. 2b (2D slice taken at $k_{\mathrm{gd,r}}=0$) the magnitude of iso-defect frequency curves decrease with increasing $m_{\mathrm{r}}$ ($>1$), as anticipated qualitatively. In addition, we notice that the iso-defect frequency curves in both Fig. 2a,b tend to asymptote to a constant value of $k_{\mathrm{gd,r}}$ and $m_{\mathrm{r}}$, respectively, at large $\mathrm{N}$. This observation provides valuable insight for our analytical model, where we can arbitrarily choose a large DMoA PnC size ($\mathrm{N}$) to obtain an accurate prediction for $\Omega_{\mathrm{d}}$, independent of the actual size of the finite DMoA PnC.
@@
Finally, Fig. 2(c-i,iii) plots the $\{\Omega_{\mathrm{d}},\kappa_{I}\}$, calculated from matrix eigen-analysis of a finite DMoA PnC ($\mathrm{N}=21$) as a function of $\{m_{\mathrm{r}},k_{\mathrm{gd,r}}\}$. A closer look at the matrix eigenvalue problem reveals a tridiagonal matrix form for the dynamic stiffness, $\mathbf{K}_{\mathrm{d}}$ (Sec. A1) that is consistent with the general form of a perturbed tridiagonal Toeplitz matrix[43, 28]. The characteristic equation for $\mathbf{K}_{\mathrm{d}}$ can be written in the implicit form:
@@
$f(\Omega)=\frac{\sin\left([n+1]\left[\cos^{-1}\left(\frac{d^{2}}{2a^{2}}-1\right)\right]\right)}{\sin\left(n\left[\cos^{-1}\left(\frac{d^{2}}{2a^{2}}-1\right)\right]\right)},$
@@
$d=2+k_{\mathrm{g,r}}-\Omega^{2},a=-1,\beta=-1,$